OpenStudy (anonymous):

What are the solutions of 6x2 − x − 15 = 0?

OpenStudy (anonymous):

Use the quadratic formula.

OpenStudy (anonymous):

\[x=(1 \pm \sqrt{1+4*6*15})/12=(1\pm19)/12\]

OpenStudy (anonymous):

huh? im confused..would that beeee 5/6..?

OpenStudy (anonymous):

(1+19)/12 =20/12=5/3 (1-19)/12 = -18/12 =-3/2

OpenStudy (anonymous):

ohh okay lol thank you :)

OpenStudy (anonymous):

yw